l=(2x+5)w=(3x-3) perimeter =64 what is length

Answer by Boreal(15235)

solve first for x
half perimeter is L+W=5x+2=32
5x=30
x=6
length is 2(6)+5=17 ANSWER
width is 3(6)-3=15
perimeter is 2(17)+2(15)=64